OpenStudy (anonymous):

Which equation can be used to find the answer? Scott works a total of 14.5 hours after school from Monday to Friday. After the weekend, he has worked a total of 27.75 hours for the entire week. How many hours does Scott work on the weekend? A. w = 14.5 + 27.75 w = 42.25 42.25 hours B. w = 14.5 + 27.75 w = 41.25 41.25 hours C. w + 14.5 = 27.75 w = 13.25 13.25 hours D. w – 14.5 = 27.75 w = 13.75 13.75 hours

OpenStudy (anonymous):

I think it's C tho

OpenStudy (anonymous):

Which equation can be used to find the answer? An auto part warehouse shipped 96 cases of oil filters and still had 172 cases on its shelves. How many cases did it have on its shelves before the shipment? A. c + 96 = 172 c = 76 76 cases B. c ÷ 2 = 96 c = 192 192 cases C. c ÷ 2 = 172 c = 344 344 cases D. c – 96 = 172 c = 268 268 cases

OpenStudy (anonymous):

how about that one

OpenStudy (anonymous):

D

OpenStudy (anonymous):

lol sorry for the late reply :D